Homework 2 11-721: Grammars and Lexicons Fall term, 2002 Due Monday, September 15

1 Based on Radford, Chapter 2, Exercise 9

Read Exercise 9 on page 108 of Radford. Consider these hypotheses: • Hypothesis 1: my coffee black is a constituent.

• Hypothesis 2: my coffee black is not a constituent.

• Hypothesis 3: the people present is a constituent.

• Hypothesis 2: the people present is not a constituent. Now do the following tasks.

• Task 1: Apply a movement test to Sentence 1.

• Task 2: Apply a movement test to Sentence 2.

• Task 3: Apply a pronominalization test to Sentence 1.

• Task 4: Apply a pronominalization test to Sentence 2.

• Task 5: Apply the coordination test to Sentence 1.

• Task 6: Apply the coordination test to Sentence 2.

• Task 7: Does the data support Hypothesis 1, Hypothesis 2, both, or neither?

• Task 8 : Does the data support Hypothesis 3, Hypothesis 4, both, or neither?

2 Based on Radford, Chapter 2, Exercise 4

Do Exercise 4 on pages 106-107 of Radford. The answers should be straightforward based on Chapter 2.